state whether the following numbers are rational or irrational . if they are irrational and expressed in the form of p/q where p and q are co prime integers then what can you say about the prime factors of q ?
a) 279.034 b)3.010010001... c)39.546782 d)2.3476817681 e) 59.120120012000...

Answer by stanbon(57347)

state whether the following numbers are rational or irrational . if they are irrational and expressed in the form of p/q where p and q are co prime integers then what can you say about the prime factors of q ?
a) 279.034
rational
= 279034/1000
-----------------------
b) x = 3.010010001...
irrational
-------
c)39.546782
rational
= 39546782/1,000,000
------------------
d) x = 2.3476817681
rational
100x = 234.76817681...
10^6x = 2347681.76817681...
(10^6-10^2)x = 2347447/999900
-------------------------------
e) x = 59.120120012000...
irrational
=======================
Cheers,
Stan H.
================